Two cars leave from the same place at the same time and travel in opposite direction.
One car travel at a constant velocity of 55 km/h and the other at 75 km/h. After how many
hours will there be 520 km apart?
1
Expert's answer
2021-07-23T16:55:02-0400

Gives

"V_1=55km/h \\
V_2=75km/h\\s=520km\\"

Both moue opposite direction

Then relatively velocity


"V=V_1+V_2=55km\/h+75Km\/h=130km\/h"

Times

"t=\frac{S}{V_1+V_2}=\frac{520}{130}=4h"